# Efficient Modelling of Anharmonicity and Quantum Effects in PdCuH2 with Machine Learning Potentials

### v1

@@ I Introduction
In the framework of standard Density Functional Theory (DFT), two fundamental approximations are commonly made. The first is concerned with the classical treatment of the nuclei, assumed to be fixed points in space generating an effective potential within which the electrons move. The second is the harmonic treatment of this potential, which is approximated through a second-order expansion in the vicinity of its local minima. These approximations are key for enabling the calculation of thermodynamic properties such as the harmonic free energy, the harmonic vibrational spectra, and the heat capacity of chemical compounds and materials [1].
In some circumstances however, these fundamental approximations are no longer valid. Two examples are systems at high temperatures whose atoms undergo large amplitude atomic vibrations, or those containing light nuclei whose quantum oscillations cannot be disregarded. In both cases the atoms explore features of the potential extending far beyond the limits of a second-order expansion around the local minima [2], defining the presence of anharmonicity, driven by temperature, or quantum nuclear effects. Anharmonicity is key for a broad range of materials properties including the emergence of charge density waves (CDW) in materials [3, 4], the enhancement or suppression of ferroelectricity in perovskites [5], and thermodynamic properties of materials at planetary temperatures [6]. However, the effects of anharmonicity are most prominent in hydrogen-rich systems, where they can alter structural phase stability, renormalize phonon spectra and dramatically modify superconducting properties [7, 8, 9, 10, 11].
@@ II Results
Figure 1: SSCHA + MLIP workflow. Flowchart for the workflow that combines the Stochastic Self-Consistent Harmonic Approximation (SSCHA) with the Machine Learning Interatomic Potential (MLIP) package for the generation of moment tensor potentials (MTPs). In this flowchart, parallelograms represent input and output files, rectangles denote individual workflow steps, and rhombi indicate logic operations. Large rounded rectangles encompass sections of the program related to the SSCHA and MLIP packages. This workflow fully integrates both programs and incorporates an on-the-fly procedure for MTP training, significantly reducing the number of required DFT calculations and automating dataset training.
It is at this step that the SSCHA+MLIP integration, based upon an active learning protocol, begins: the workflow is depicted schematically in Figure 1. Displaced configurations for the supercell are created, and single point DFT calculations are performed to obtain the EFS of all of the members of this first population, to be used for training the MTP. In subsequent populations, the extrapolation grade, $\gamma$, of each structure is determined via the generalised D-optimality criterion, which is able to detect the most diverse configurations  [37]. To ensure accurate predictions we use a very conservative value of $\gamma_{\text{select}}=2$, chosen so that every geometry with $\gamma_{\text{select}}>2$, which is not well described by the MTP, has their structures and EFS added to the training set for retraining the potential.
@@
The SSCHA + MLIP workflow, including the active learning protocol and the upscaling procedure, is more efficient and complete than recently proposed workflows [5, 42]. A key advantage of our approach is that it enables the automation of the training and facilitates the evaluation of its efficacy, ensuring accuracy in the predictions, via the active learning phase. Additionally, this methodology not only provides the possibility of drastically decreasing the computational cost required for calculating the EFS during the upscaling procedure, but it also enables the identification of emerging phenomena such as CDWs, which potentially require retraining of the potential. Furthermore, a MLP developed at a particular temperature and pressure can be employed as a starting guess for a simulation performed for a different set of thermodynamic variables. For example, an MTP trained at high temperatures could be used for a calculation performed at low temperatures, requiring only a few additional DFT calculations for retraining via the active learning protocol. Similarly, an MTP trained at a few pressures in a narrow pressure range could be employed to perform the SSCHA dynamics allowing for a full cell relaxation with minimal computational expense.
Figure 3: Crystal structures. The most stable PdCuH2 phases proposed in (a) Ref. [46], and (b) that found here. Both studies identified the same PdCuH phase as the ground state. It can be obtained from $P4/mmm$ PdCuH2 by removing the hydrogen atoms that are at the same sites as the hydrogen in the center of the blue transparent octahedron.
@@
The phonon band structure for the lower symmetry $I4_{1}/amd$ PdCuH2 phase is markedly different. Importantly, no instabilities are present in the harmonic phonons and the introduction of quantum nuclear effects does not significantly affect the vibrational frequencies associated with a particular mode, suggesting the motions of atoms in this phase are mostly harmonic. The most notable difference between the harmonic and the SSCHA-computed phonon modes are the minor renormalizations of the intermediate and upper optical branches, which undergo a blue shift separating them in energy.
@@
Considering the similarity of PdCuH2 with PdH, we wondered if it might be subject to the inverse isotope effect? Towards this end, we calculated the phonon spectrum, Eliashberg spectral function (Supplementary Figure 2) and resulting $T_{\text{c}}$ of the deuterated compound with and without the SSCHA. While the $\lambda$ of $P4/mmm$ PdCuH2 and PdCuD2 were nearly identical, the latter had a reduction of $\sim$15in $\omega_{\textrm{log}}$, as expected, resulting in a decreased $T_{\text{c}}$ of 11-17 K. Therefore, no inverse isotopic effect was found.
@@ III Discussion
This study outlines an effective protocol that combines the Stochastic Self-Consistent Harmonic Approximation (SSCHA) with the Machine Learning Interatomic Potential (MLIP) package, enabling the automated and rapid calculation of anharmonic and quantum nuclear effects for a wide range of materials properties. In the proposed workflow DFT calculations are performed on geometries (generated by the Monte Carlo algorithm of the SSCHA) that are deemed to be extrapolative, and a subset of these is employed to train quantum-accurate moment tensor potentials (MTPs) that are used to predict the energies, forces and stresses of the remaining geometries required for the SSCHA. Furthermore, we introduce an upscaling protocol, where the MTPs are first trained on small cells, and retrained on larger ones, allowing for the simulation of systems with sizes that were previously out-of-computational-reach. The upscaling protocol is based on the assumption that most of the interatomic interactions are already captured in the small cells, and that only minor corrections to the MTPs are required when larger cells are considered. This procedure opens the door for the routine modelling of quantum and anharmonic effects, on cells of up to thousands of atoms, while at the same time dramatically enhancing the computational efficiency.
@@
The protocol is put to the test on the PdCuH2 system, which was hypothesized [46] to be an important contributor to the superconductivity reported for hydrogenated PdCu alloys with a maximum superconducting critical temperature, $T_{\text{c}}$, of 17 K at ambient pressures [45]. Hydrogenation of the most stable $Pm\bar{3}m$ symmetry PdCu phase would likely yield the kinetically favored $P4/mmm$ PdCuH2 structure, metastable by 18 meV/atom from the newly predicted $I4_{1}/amd$ PdCuH2 ground state. Similar to the highly anharmonic PdH phase, $P4/mmm$ PdCuH2 is found to be dynamically stable only upon the inclusion of quantum and anharmonic effects, which heal the imaginary modes in its phonon band structure, and renormalize the frequencies of others. These effects decrease the predicted $T_{\text{c}}$ of $P4/mmm$ PdCuH2 from 37-44 K to 13-20 K, mirroring previous trends computed for PdH [52], whose anharmonic $T_{\text{c}}$ is notably smaller ($\sim$5 K), likely because of the lighter mass of the copper, as compared to the palladium, atoms. An inverse isotope effect on superconductivity was not observed when hydrogen was replaced by deuterium, with an estimated $T_{\text{c}}$ of 11-17 K for $P4/mmm$ PdCuD2. The thermodynamically preferred $I4_{1}/amd$ PdCuH2 phase was shown to be relatively well treated within the harmonic approximation, with the predicted $T_{\text{c}}$ decreasing from 5-10 K to 2-5 K upon inclusion of anharmonicity.
@@
The good agreement between the computed $T_{\text{c}}$ of $P4/mmm$ PdCuH2 and that measured for the hydrogenated alloy, with composition likely being closer to Pd1.1Cu0.9H1.4, suggests structural similarities, on a local level. Further, the results presented herein may have broader implications for the discovery of ambient-condition superconducting hydrides, as many of the recently predicted compounds involve ternaries featuring Pd (e.g. Mg2PdH6 [53]), Cu (e.g. Li2CuH6 [54]) and other combinations of alkaline earth and transition metal atoms [55]. In addition This procedure opens the door towards studies of anharmonicity using more costly electronic structure methods, such as hybrid or meta-GGA functionals, which were previously out of reach. Furthermore, it enables a determination of the relative stability of enthalpically close phases, which could not be separated well by the PBE functional as for example in the high pressure phases of lithium [56].
@@ IV.1 Computational Details
The CSP searches were performed using the XtalOpt open-source evolutionary algorithm [57, 58, 59, 60]. The initial generation consisted of random symmetric structures created by randspg [61], and duplicate structures were identified via XtalComp [61] and discarded from the breeding pool. The DFT calculations for the CSP were performed with the Vienna ab-initio Simulation Package (VASP) version 5.4.1 [62, 63], while the calculations of the electronic band structures, relative enthalpies, and the SSCHA were performed with the Quantum ESPRESSO (QE) [64, 65] package. PAW and ultrasoft pseudopotentials were used for VASP and QE, respectively, both within the Perdew-Burke-Ernzerhof (PBE) [66] parametrization of the exchange-correlation functional. The Pd 4d95s15p04f0, Cu 3d104s14p04f0, H 1s12p0, and Pd 5s15p04d9, Cu 4s1.54p03d9.5, H 1s1 electronic configurations were treated explicitly for VASP and QE, respectively. The CSP relaxations for each structure were performed using a 4 step procedure where (i) the atomic positions, (ii) cell parameters and (iii+iv) cell parameters and ionic positions were relaxed. In these steps the energy cutoffs for the wavefunctions and a $\mathbf{k}$-point grid density were set to 120, 240, 300 and 400 eV, and $2\pi\times 0.1$, $2\pi\times 0.07$, $2\pi\times 0.05$, and $2\pi\times 0.025$1, respectively. The single point QE calculations were performed imposing energy cutoffs for the wavefunctions and density as 1224 eV and 12,240 eV, respectively. The integration over the Brillouin zone in the self-consistent calculations was performed with a first-order Methfessel-Paxton [67] smearing of 0.272 eV coupled with a $\mathbf{k}$-point grid with density $2\pi\times 0.0175$1.
@@
Additional Information: Supplementary information The online version contains supplementary material available at

### v2

@@ I Introduction
In the framework of standard Density Functional Theory (DFT), two fundamental approximations are commonly made. The first is concerned with the classical treatment of the nuclei, assumed to be fixed points in space generating an effective potential within which the electrons move. The second is the harmonic treatment of this potential while calculating the vibrational properties, which is approximated through a second-order expansion in the vicinity of its local minima. These approximations are key for enabling the calculation of thermodynamic properties such as the harmonic free energy, the harmonic vibrational spectra, and the heat capacity of chemical compounds and materials [1].
However, in some circumstances these fundamental approximations are no longer valid. Two examples are systems at high temperatures whose atoms undergo large amplitude atomic vibrations, or those containing light nuclei whose quantum oscillations cannot be disregarded. In both cases the atoms explore features of the potential extending far beyond the limits of a second-order expansion around the local minima [2], defining the presence of anharmonicity, driven by temperature, or quantum nuclear effects. Anharmonicity is key for a broad range of materials properties including the emergence of charge density waves (CDW) in materials [3, 4], the enhancement or suppression of ferroelectricity in perovskites [5], and thermodynamic properties of materials at planetary temperatures [6]. The effects of anharmonicity are most prominent in hydrogen-rich systems, where they can alter structural phase stability, renormalize phonon spectra and dramatically modify superconducting properties [7, 8, 9, 10, 11].
@@ II Results
Figure 1: SSCHA + MLIP workflow. Flowchart for the workflow that combines the Stochastic Self-Consistent Harmonic Approximation (SSCHA) with the Machine Learning Interatomic Potential (MLIP) package for the generation of moment tensor potentials (MTPs). In this flowchart, the blue rectangles denote individual workflow steps, the orange rhombi indicate logic operations, while the generic yellow parallelograms represent input and output files. Large rounded rectangles encompass sections of the program related to the SSCHA and MLIP packages. This workflow fully integrates both programs and incorporates an on-the-fly procedure for MTP training, significantly reducing the number of required DFT calculations and automating dataset training.
It is at this step that the SSCHA+MLIP integration, based upon an active learning protocol, begins: the workflow is depicted schematically in Figure 1. Displaced configurations for the supercell are created, and single point DFT calculations are performed to obtain the EFS of all of the members of this first population, to be used for training the MTP. In subsequent populations, the extrapolation grade, $\gamma$, of each structure is determined via the generalised D-optimality criterion, which is able to detect the most diverse configurations  [37]. To ensure accurate predictions we use a very conservative value of $\gamma_{\text{select}}$ = 2, chosen so that every geometry with $\gamma_{\text{select}}>$ 2 has their structures and EFS added to the training set. This provides a strict threshold for the selection of the training set.
@@
The SSCHA + MLIP workflow, including the active learning protocol and the upscaling procedure, is more efficient and complete than recently proposed workflows [5, 42]. A key advantage of our approach is that it enables the automation of the training and facilitates the evaluation of its efficacy, ensuring accuracy in the predictions, via the active learning phase. The previously established methods require prior knowledge of the final structure and its environment to effectively train the potential. In contrast, the active learning protocol proposed here enables the use of an arbitrary initial structure. Additionally, the accuracy of this model was evaluated by comparing its errors to those reported by Lucrezi et al. [42]. For a 2$\times$2$\times$2 supercell and a level 10 potential, our MTP errors are 0.5-0.6 meV/atom for the energies, 20-30 meV/Å for the force components, and 0.2-0.3 GPa for the diagonal stress tensor components. These results are in good agreement with those presented by Lucrezi et al. [42]. More details on the errors obtained for the potential training are reported in Supplementary Note 1.
Additionally, this methodology not only provides the possibility of drastically decreasing the computational cost required for calculating the EFS during the upscaling procedure, but it also enables the identification of emerging phenomena such as CDWs, which potentially require retraining of the potential. Furthermore, a MLP developed at a particular temperature and pressure can be employed as a starting guess for a simulation performed for a different set of thermodynamic variables. For example, an MTP trained at high temperatures could be used for a calculation performed at low temperatures, requiring only a few additional DFT calculations for retraining via the active learning protocol. Similarly, an MTP trained at a few pressures in a narrow pressure range could be employed to perform the SSCHA dynamics allowing for a full cell relaxation with minimal computational expense.
Figure 3: Crystal structures. The most stable PdCuH2 phases proposed in (a) Ref. [46], and (b) that was found here. Both studies identified the same PdCuH phase as the ground state. It can be obtained from $P4/mmm$ PdCuH2 by removing the hydrogen atoms that are at the same sites as the hydrogen in the center of the blue transparent octahedron.
@@
A final relaxation of the crystal structure was performed through the SSCHA at zero temperature for both the $P4/mmm$ and the $I4_{1}/amd$ phases after performing CSP. In this relaxation the unit cell shape and volume were held fixed, while allowing the atomic positions to relax, subject to symmetry constraints. No significant modifications to the crystal structure were observed as a result of this relaxation. Furthermore, the introduction of quantum nuclear effects did not induce any anisotropic pressure contributions, eliminating the necessity for a further relaxation of the unit cell.
@@
To ensure that the anharmonic corrections do not derive from artifacts of the local properties of the potential we used the meta-GGA R2SCAN functional, as implemented in the VASP code, to calculate both the harmonic and quantum anharmonic phonons (Supplementary Figures 6 and 7). Mirroring the results obtained with the PBE functional and the Quantum Espresso code, the SSCHA healed an imaginary mode at the $Z$-point, as well as soft modes around the $R$ and $X$ points, rendering this phase dynamically stable. Furthermore, the comparison between the quantum anharmonic phonons obtained with PBE and R2SCAN (Supplementary Figure 6) suggests that the R2SCAN introduces just minor corrections to the spectra, confirming that the PBE provides a good approximation for the system.
The phonon band structure for the lower symmetry $I4_{1}/amd$ PdCuH2 phase is markedly different. Importantly, no instabilities are present in the harmonic phonons and the introduction of quantum nuclear effects does not significantly affect the vibrational frequencies associated with a particular mode, suggesting the motions of atoms in this phase are mostly harmonic. The most notable difference between the harmonic and the SSCHA-computed phonon modes are the minor re-normalisation of the intermediate and upper optical branches, which undergo a blue shift separating them in energy.
@@
Considering the similarity of PdCuH2 with PdH, the possible presence of the inverse isotope effect was investigated. Towards this end, the phonon spectrum, Eliashberg spectral function (Supplementary Figure 4) and the resulting $T_{\text{c}}$ of the deuterated compound was calculated with and without the SSCHA. While the $\lambda$ of $P4/mmm$ PdCuH2 and PdCuD2 were nearly identical, the latter had a reduction of $\sim$15in $\omega_{\textrm{log}}$, as expected, resulting in a decreased $T_{\text{c}}$ of 11-17 K. Therefore, no inverse isotopic effect was found.
@@
The renormalization of the $T_{\mathrm{c}}$ in both structures can be predominantly linked to the change of the phonon spectra. Lucrezi et al. [42] proposed that the discrepancy in the phonon spectra between the SSCHA and the harmonic calculations can be separated into two contributions: structural renormalizations induced by quantum nuclear effects and phonon-phonon interactions driven by anharmonicity. For the PdCuH2 structures, since no structural renormalization arises from SSCHA, the phonon renormalization is solely a consequence of anharmonicity-induced phonon-phonon interactions.
@@ III Discussion
This study outlines an effective protocol that combines the Stochastic Self-Consistent Harmonic Approximation (SSCHA) with the Machine Learning Interatomic Potential (MLIP) package, enabling the automated and rapid calculation of anharmonic and quantum nuclear effects for a wide range of materials properties. This procedure enhances the performance of the SSCHA, improving the prospect for a routine treatment of quantum nuclear and anharmonic effects in computations. In the proposed workflow DFT calculations are performed on geometries (generated by the Monte Carlo algorithm of the SSCHA) that are deemed to be extrapolative, and a subset of these is employed to train quantum-accurate moment tensor potentials (MTPs) that are used to predict the energies, forces and stresses of the remaining geometries required for the SSCHA. Furthermore, we introduce an upscaling protocol, where the MTPs are first trained on small cells, and retrained on larger ones, allowing for the simulation of systems with sizes that were previously out-of-computational-reach. The upscaling protocol is based on the assumption that most of the interatomic interactions are already captured in the small cells, and that only minor corrections to the MTPs are required when larger cells are considered.
Although general in nature, this procedure is not directly applicable to other types of MLPs due to the specific implementation of the active learning protocol in the MLIP program package. An extension of the procedure to GAPs, could be possible by adopting an active learning strategy similar to the one proposed in Refs.  [53, 54], where new structures for the training set are selected on the estimation of the Bayesian error during the simulation. Alternatively, for neural networks, the active learning procedure could involve randomly selecting structures to create a validation set at each step of the simulation, which would then be used to assess the accuracy of the potential.
@@
The protocol is put to the test on the PdCuH2 system, which was hypothesized [46] to be an important contributor to the superconductivity reported for hydrogenated PdCu alloys with a maximum superconducting critical temperature, $T_{\text{c}}$, of 17 K at ambient pressure [45]. Hydrogenation of the most stable $Pm\bar{3}m$ symmetry PdCu phase would likely yield the kinetically favored $P4/mmm$ PdCuH2 structure, metastable by 18 meV/atom from the newly predicted $I4_{1}/amd$ PdCuH2 ground state. Similar to the highly anharmonic PdH phase, $P4/mmm$ PdCuH2 is found to be dynamically stable only upon the inclusion of quantum and anharmonic effects, which heal the imaginary modes in its phonon band structure, and renormalize the frequencies of others. These effects decrease the predicted $T_{\text{c}}$ of $P4/mmm$ PdCuH2 from 37-44 K to 13-20 K, mirroring previous trends computed for PdH [52], whose anharmonic $T_{\text{c}}$ is notably smaller ($\sim$5 K), likely because of the lighter mass of the copper, as compared to that of palladium. An inverse isotope effect on superconductivity was not observed when hydrogen was replaced by deuterium, with an estimated $T_{\text{c}}$ of 11-17 K for $P4/mmm$ PdCuD2. The thermodynamically preferred $I4_{1}/amd$ PdCuH2 phase was shown to be relatively well treated within the harmonic approximation, with the predicted $T_{\text{c}}$ decreasing from 5-10 K to 2-5 K upon inclusion of anharmonicity.
@@
The good agreement between the computed $T_{\text{c}}$ of $P4/mmm$ PdCuH2 and that measured for the hydrogenated alloy, with composition likely being closer to Pd1.1Cu0.9H1.4, suggests structural similarities, on a local level. Further, the results presented herein may have broader implications for the discovery of ambient-condition superconducting hydrides, as many of the recently predicted compounds involve ternaries featuring Pd (e.g. Mg2PdH6 [55]), Cu (e.g. Li2CuH6 [56]) and other combinations of alkaline earth and transition metal atoms [57]. In addition this procedure opens the door towards studies of anharmonicity using more costly electronic structure methods, such as hybrid or meta-GGA functionals, which were previously out of reach. This enables a determination of the relative stability of enthalpically close phases, which could not be separated well by the PBE functional as for example in the high pressure phases of lithium [58].
To conclulde we also mention the limitations of this method. Due to the constraitn related to the Gaussian density matrix, the SSCHA is inadequate for describing molecular solids with both librational and free rotational modes [59]. Furthermore, it fails to capture tunnelling states and struggles when the probability distribution deviates significantly from a Gaussian form [60]. However this technique performs well for atomic and ionic solids, where the point-like nature of the atomic or ionic building blocks suppresses large-amplitude curvilinear motion. For such systems, the SSCHA offers a significantly more computationally efficient alternative compared to Path Integral Molecular Dynamics (PIMD) [59].
@@ IV.1 Computational Details
The CSP searches were performed using the XtalOpt open-source evolutionary algorithm [61, 62, 63, 64]. The initial generation consisted of random symmetric structures created by randspg [65], and duplicate structures were identified via XtalComp [65] and discarded from the breeding pool. The DFT calculations for the CSP were performed with the Vienna ab-initio Simulation Package (VASP) version 5.4.1 [66, 67], while the calculations of the electronic band structures, relative enthalpies, and the SSCHA were performed with the Quantum ESPRESSO (QE) [68, 69] package (unless otherwise noted). PAW and ultrasoft pseudopotentials were used for VASP and QE, respectively, both within the Perdew-Burke-Ernzerhof (PBE) [70] and R2SCAN meta-GGA [71] parametrization of the exchange-correlation functional. The Pd 4d95s15p04f0, Cu 3d104s14p04f0, H 1s12p0, and Pd 5s15p04d9, Cu 4s1.54p03d9.5, H 1s1 electronic configurations were treated explicitly for VASP and QE, respectively. The CSP relaxations for each structure were performed using a 4 step procedure where (i) the atomic positions, (ii) cell parameters and (iii+iv) cell parameters and ionic positions were relaxed. In these steps the energy cutoffs for the wavefunctions and a $\mathbf{k}$-point grid density were set to 120, 240, 300 and 400 eV, and $2\pi\times 0.1$, $2\pi\times 0.07$, $2\pi\times 0.05$, and $2\pi\times 0.025$1, respectively. The single point QE calculations were performed imposing energy cutoffs for the wavefunctions and density as 1224 eV and 12,240 eV, respectively. The integration over the Brillouin zone in the self-consistent calculations was performed with a first-order Methfessel-Paxton [72] smearing of 0.272 eV coupled with a $\mathbf{k}$-point grid with density $2\pi\times 0.0175$1.
@@
The anharmonic $T_{\mathrm{c}}$s were calculated using the phonon eigenvalues and eigenvectors obtained through the SSCHA relaxation, along with the electron-phonon coupling matrix elements obtained from the classical DFPT calculations for the same crystal structure, as the SSCHA relaxation did not result in any significant structural changes.